without graphing, determine which of the following 3 points:
P1 = (8,6)
P2 = (2,5)
P3 = (4,1)
are part of the graph of the following system:
y - 10x <= 0
2y - 3x >= 0
y + x <= 15

Answer by solver91311(24713)

without graphing, determine which of the following 3 points:
P1 = (8,6)
P2 = (2,5)
P3 = (4,1)
are part of the graph of the following system:
y - 10x <= 0
2y - 3x >= 0
y + x <= 15

Just substitute the values for the x- and y-coordinates into the inequalities. If all three statements are true for a point, then the point is in the solution set, but if any one of the statements is false, the point is not in the solution set.
6 - 10(8) <= 0: True
2(6)-3(8) >= 0: False. Don't bother checking the third inequality, the point P1(8,6) is not in the solution set.

Follow the same procedure for the other two points.